OpenStudy (anonymous):

If g(3) = 12, then g(x) = 2x + _____?

OpenStudy (anonymous):

+6

OpenStudy (anonymous):

let g(x) = 2x + k g(3)= 2(3)+k 12 = 6+k k=6